OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4

OpenStudy (johnweldon1993):

you can use the equaation \[\large y - y_0 = m(x - x_0)\] you plug in you slope for 'm' ...your x coordinate for Xo and your y coordinate for Yo

OpenStudy (anonymous):

how do I know what else to put for the x1, and y1?

OpenStudy (johnweldon1993):

What do you mean? well you leave the 'y' and the 'x' alone...those are part of your final equation...the only thing you worry about in that equation are 'm' 'Xo' and 'Yo' you have slope (4) this is your 'm' you have the point (-3 , -1) these are you x and y coordinates....just replace everything

OpenStudy (anonymous):

I still don't understand.. I tried plugging in the numbers into the equation, and the answer I got doesn't match any of the answers I have available.

OpenStudy (johnweldon1993):

So what I get is... \[\large y - y_0 = m(x - x_0)\] after replacing everything with the numbers you have I get: \[\large y - (-1) = 4(x - (-3))\] so lets simplify that a bit... \[\large y + 1 = 4(x + 3)\] now lets distribute out that 4... \[\large y + 1 = 4x + 12\] And finally subtract 1 from both sides... \[\large y = 4x + 11\]

OpenStudy (johnweldon1993):

Does that make more sense @jocelynsalcedo ?

OpenStudy (anonymous):

my problem was I didn't distribute. thanks.

OpenStudy (johnweldon1993):

Ahh okay....glad you understand now.....and anytme! :)